OpenStudy (anonymous):

What is the value of the function y = 0.5x – 10 when x = 9?

OpenStudy (gorv):

plug in x=9

OpenStudy (gorv):

y=0.5*9-10

OpenStudy (gorv):

what u got??

OpenStudy (anonymous):

Its -5.5 right?

OpenStudy (gorv):

0.5*9=4.5

OpenStudy (gorv):

4.5-10=-5.5

OpenStudy (gorv):

well done :)

OpenStudy (anonymous):

5.5